40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10512     Entry time: Wed Sep 17 01:40:55 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Closer to REFL DC? Maybe? 

I tried a bunch of times to reduce my CARM offset so I could jump to REFLDC digitally, but I think I'm maybe being a little ambitious with the arm power I'm trying to get to.

I have modified the carm_cm_up script so that it does my new procedure.  Everything is the same through locking the PRCL and MICH on 3f.  Then it reduces the CARM offset to 1.5 nm.  This is where we *used* to transition to sqrtInvTrans.  Now I have it going a bit farther to 0.5 nm, and arm powers of about 1 before doing that transition.  Also, before it transitions it lowers the CARM gain and engages the 1kHz lowpass in FM9.  A gain of about 4 is fine to keep the gain peaking in the CARM loop to only about 10dB, and sets a UGF of 100Hz which is the peak of the phase bubble with the lowpass engaged. 

Once I got to this point (several times tonight), I turned on CARM and DARM oscillations and looked at the transfer functions between (CARM and REFLDC) and (DARM and AS55Q). I have 2 DTT templates setup for this, in /users/Templates/PRFPMI.  These templates assume that you have your new DARM signal (AS55) going to SRCL_IN1 and your new CARM signal (REFLDC, which is actually REFL11I coming through the CM board) going to MC_IN1. 

I'm not sure why I'm losing lock. I don't see anything terribly telling on the time series plots, in particular none of the loops look like they are oscillating.  Here is one of the better examples from this evening:

Zoom_TRXTRYupto25_1094967592.png

Other notes:

* I realigned the Xgreen on the PSL table (again) to maximize the beatnote amplitude.  Y was fine, but X was very poorly overlapped on the camera.

* I put the SR785 back by the LSC rack and plugged it into the CM board for transfer functions.  Didn't take any tonight.

* We have a small wishlist for scripting things:  (1) DRMI restore script should reset REFL11 to "normal" REFL11.  (2) CARM/DARM acquisition restore script should reset REFL11 to REFLDC.  (3) CARM/DARM acquisition restore should also set PRMI parameters (as Q noted last week).

ELOG V3.1.3-